OpenStudy (anonymous):

Compare the balance after 25 years of a $10,000 investment earning 6.75% interest compounded continuously to the same investment compounded semiannually.

OpenStudy (anonymous):

\[10,000\times e^{.0675\times 25}\] vs \[10,000\times (1+\frac{.0675}{2})^{50}\] then a calulator

OpenStudy (anonymous):

So, the answer would be 52574.61826?

OpenStudy (anonymous):

i don't know i didn't do it. but you have to compare the two answers

OpenStudy (anonymous):

The first equation gave me 5232.481438 The second equation gave me 52574.61826

OpenStudy (anonymous):

hmm then something is wrong because first one should be larger than second one

OpenStudy (anonymous):

I don't know what to do now? Are those the answers?